# Muchisim: A Simulation Framework for Design Exploration of Multi-Chip Manycore Systems

### v1

@@ IV-A Validation
We compare the runtimes reported on the WSE for Fast Fourier Transforms (FFTs) [65] with the runtimes obtained with our simulator. We compare that with all their reported datapoints, i.e., parallelizing FFT of $n^{3}$ tensors across $n^{2}$ processors, for $n$ equal to 32, 64, 128, 256, and 512. In doing so, we show that Muchisim has high level of scalability and the accuracy is not impacted by the size of the DUT.
The WSE mounted on the CS-2 [51] is a 7nm-based 46,${}^{2}$ monolithic die with 850,000 cores and 40GB of SRAM. To model the WSE, we configure the DUT as a package with a single chiplet (32-bit 2D mesh NoC) with no integrated DRAM [51]. The WSE has a circuit-switched NoC with traffic filtering, while our simulator models a packet-switched NoC. The advantage of the circuit-switched NoC is to avoid message headers and simplify the routers, but it requires more synchronization when setting up a path to send data. To overcome this difference, we do not consider the destination header in this case, and rather consider the same amount of data sent by the PUs in both cases. The performance model for the PUs is set based on their reported numbers [65].
Runtime: The runtimes of the $n^{3}$ FFT problems parallelized across $n^{2}$ tiles reported on the WSE [65] are 1.2×1.2$\times$1.2 × of the runtimes reported by Muchisim, consistently, for $n$ ranging from 32 to 512 (i.e., from a thousand to a quarter million tiles). We believe the reason for this slightly optimistic runtime is that Muchisim does not model the synchronization overhead of the circuit-switched NoC. We also simulated these datapoints with up to 32 host threads to measure the speedup of the simulator, shown in fig. 3 for FFT and other applications.
In terms of chip area, the area reported by the simulator is 8.8% larger than the area of the WSE. We attribute this to WSE’s NoC, where their routers are likely smaller than the default area of Muchisim’s routers. Regarding energy, the simulator reports an average power of the processing tiles of a bit over 1KW. While Morenes et. al.  [65] did not report energy for the FFT evaluation, we know that the entire CS-2 system (including cooling) draws up to 20KW [51]. Considering that 512x512 is less than a third of the CS-2, and that PU utilization is low ($\sim$30%) because of the communication bottleneck of FFT, the energy estimation seems reasonable.
@@ IV-B Parallelization Speedup and Scaling Throughput
Figure 3: Ratio between the simulator and the DUT runtime, for two DUT sizes (32x32 and 64x64 tiles, monolithic, connected via a 64-bit Torus), evaluated with an increasing number of host threads to process the same R22 dataset. The DUT time is considered as the aggregated runtime of all tiles. The simulator runtime decreases close to linearly with the number of host threads.
@@
We show the speedup of Muchisim with the number of host threads in fig. 3, for two DUT sizes, as the ratio between the simulator and the DUT runtime. The ratio of the host’s simulation time to the simulated runtime of executing these applications on the DUT decreases from a geomean ratio of 614 to 43 (a 12$\times$ speedup) across the datapoints, when scaling from 2 to 32 host threads (Intel Xeon Gold 6342 at 2.8Ghz). The speedup is Muchisim’s parallelization is perfectly linear until the point where each host thread only processes a couple of tile columns (32-thread 32x32 DUT). fig. 3 demonstrates the efficiency of Muchisim and its parallelization. On a 32-thread host, simulating the execution of applications on the DUT takes only 43×43$\times$43 × (on geomean, and down to 21×21$\times$21 × for FFT) longer than the expected runtime of every tile of the DUT. This ratio is remarkable considering that Muchisim performs a functional simulation of data-dependent workloads, where all NoC routers are evaluated every cycle at the flit level.
@@
Figure 4: Simulation time (in host seconds) and throughput in DUT operations and NoC message fits routed per second, for scaling DUT sizes from a thousand to a million tiles (32x32 tiles per chiplet, connected via a 64-bit hierarchical 2D Torus) when processing the R26 dataset. The $2^{10}$ and $2^{12}$ datapoints are evaluated with 16 and 32 host threads, respectively, of a single-socket Intel Xeon Gold 6342 at 2.8Ghz, the datapoints from $2^{14}$ to $2^{18}$ with 64 and $2^{20}$ with 128 threads, respectively, of a 4-socket Intel Xeon Gold 6230 at 2.1GHz.
@@
fig. 4 shows the simulation time and throughput of Muchisim (in host seconds) when using increasingly large DUT configurations (from a thousand to a million tiles) to execute the R26 dataset. The first three datapoints see a decreased or nearly constant simulation time because we are increasing the parallelization of Muchisim from 16 to 64 host threads. The NoC simulation throughput ranges from a few million message flits routed per second (PageRank) to over 40 million (SSSP). Note that this NoC throughput does not count the attemps to routed that fail due to contention (two input ports compete for the same output) or backpressure (the buffer of the output port is full). The throughput of operations—natively executed on the host—reaches up to a few billion.
@@
The absolute runtimes (fig. 4, blue) show that even the simulations of million-tile DUT configurations on a billion-element dataset—unattainable before Muchisim—are simulated in under half a day for BFS, SPMV and Histo, and in up to two days for SSSP and PageRank, on a single server-class chip. FFT (not shown in fig. 4 since it does not use a graph but a $n^{3}$ tensor for $n^{2}$ parallelization) is simulated under 100 seconds for up to $n=2^{7}$, i.e, $2^{14}$ tiles, but it also takes a day to simulate a million tiles on a billion-element ($1024^{3}$) tensor, where the all-to-all communication of FFT dominates the runtime.
@@ IV-C Case Study: Memory Integration
Figure 5: Performance, energy efficiency, and performance per dollar improvements of the DUT using different SRAM sizes and number of tiles per HBM channel, over a baseline of 64KB SRAM and 128 tiles per HBM channel (T/C). In this study, a chiplet is always attached to a single 8-channel HBM device, and thus, the number of tiles per chiplet (16x16 or 32x32) determines the ratio of tiles per HBM channel. The R25 dataset is studied on a DUT with 1024 tiles; the dataset footprint per tile ranges from 4MB (Histogram) to 8MB per tile (SPMV).
@@
As a case study, we studied the performance, energy efficiency, and performance per dollar improvements of different SRAM sizes and numbers of tiles per chiplet (same number of tiles in total) and show the results in fig. 5. Since all applications but SPMM have a very low arithmetic intensity with respect to the data movement, they require a large amount of memory bandwidth. We observe, (fig. 5 top), a strong performance increase with SRAM size, 3.5×3.5$\times$3.5 × on geomean when increasing from 64KB to 256KB, with the same chiplet configuration of 32x32 tiles, and an additional 2×2$\times$2 × increase when using 16x16-tile chiplets (32T/C), for the same SRAM size of 256KB. The hit-rate of the data cache (not shown) increases from 83% to 95% on geomean with the SRAM size increase, but since the effective bandwidth of a tile is $SRAM\_bandwidth\times hit\_rate+DRAM\_bandwidth\times(1-hit\_rate)$, the hit-rate has a larger impact on the bandwidth when more tiles share the same DRAM channel (128T/C).
@@
The throughput per dollar is lower with 16x16-tile chiplets for all applications but SSSP and SPMV, due to the extra cost of having four times more HBM devices. SPMM, with over an order of magnitude more arithmetic intensity than the rest of the applications for the dataset evaluated (K25), showcases that when targetting performance per dollar, the optimal memory integration depends on the application domain.
V Conclusion
@@
Muchisim allows evaluating tiled manycore architectures with different hierarchical organizations, network topologies, and memory integrations. It also supports various parallelization strategies (do-all and task-based) and communication primitives (e.g., message-passing and reduction trees). Moreover, Muchisim enables exploring the optimal balance between resources dedicated to compute, memory and network, for different target metrics and application domains. Muchisim comes with a benchmark of eight memory- and communication-intensive applications and ten datasets of several sizes and characteristics. Muchisim also includes two data visualization tools to facilitate the analysis of the simulation results.
@@
We have created the scripts for the artifact evaluation of the experiments presented in this paper. Our repository[67] includes the Muchisim simulation framework, applications, datasets and visualization tools; we are planning to document these further with the publication of this paper to allow researchers to explore the design space of scale-out manycore architecture as well as extend it with their own novel techniques. Future work on the simulator itself includes multi-node MPI parallelization and support for low-diameter cluster interconnects, to enable simulating even larger systems.

### v2

@@ IV-A Validation
We compare the runtimes reported on the WSE for Fast Fourier Transforms (FFTs) [66] with the runtimes obtained with our simulator. We perform simulations for all their reported datapoints, i.e., parallelizing FFT of $n^{3}$ tensors across $n^{2}$ processors, for $n$ equal to 32, 64, 128, 256, and 512. In doing so, we show that MuchiSim has a high level of scalability and the accuracy is not impacted by the size of the DUT.
The WSE mounted on the CS-2 [51] is a 46,225mm2 monolithic die with 850,000 cores and 40GB of SRAM on 7 nm technology. To model the WSE, we configure the DUT as a package with a single chiplet (32-bit 2D mesh NoC) with no integrated DRAM [51]. The WSE has a circuit-switched NoC with traffic filtering, while our simulator models a packet-switched NoC. The advantage of the circuit-switched NoC is to avoid message headers and simplify the routers, but it requires more synchronization when setting up a path to send data. To overcome this difference, we do not consider the destination header in this case, and rather consider the same amount of data sent by the PUs in both cases. The performance model for the PUs is set based on their reported numbers [66].
Runtime: The runtimes of the $n^{3}$ Fast Fourier Transform (FFT) problems parallelized across $n^{2}$ tiles reported on the WSE [66] are 1.2×1.2$\times$1.2 × of the runtimes reported by MuchiSim, consistently, for $n$ ranging from 32 to 512 (i.e., from a thousand to a quarter million tiles). We believe the reason for this slightly optimistic runtime is that MuchiSim does not model the synchronization overhead of the circuit-switched NoC of the WSE. We also simulated these datapoints with up to 32 host threads to measure the speedup of the simulator, shown in fig. 3 for FFT and other applications.
In terms of chip area, the area reported by the simulator is 8.8% larger than the area of the WSE. We attribute this to WSE’s NoC, where their routers are likely smaller than the default area of MuchiSim’s routers. Regarding energy, the simulator reports an average power of the processing tiles of a bit over 1 KW. While the WSE study did not report energy for the FFT evaluation, we know that the entire CS-2 system (including cooling) draws up to 20 KW [51]. Considering that 512x512 is less than a third of the CS-2, and that PU utilization is low ($\sim$30%) because of the communication bottleneck of FFT, the energy estimation seems reasonable.
@@ IV-B Parallelization Speedup and Scaling Throughput
Figure 3: Ratio between the simulator and the DUT runtime, for two DUT sizes (32x32 and 64x64 tiles, monolithic, connected via a 64-bit 2D torus), evaluated with an increasing number of host threads to process the same RMAT-22 dataset. The DUT time is considered as the aggregated runtime of all tiles. The simulator runtime decreases close to linearly with the number of host threads.
@@
fig. 3 shows the performance of MuchiSim with the number of host threads, for two DUT sizes. This is shown as the ratio between the host runtime for the simulator and the simulated runtime of executing these applications on the DUT runtime. This ratio decreases from a geomean ratio of 614 to 43 (a 12$\times$ speedup) across the datapoints, when scaling from 2 to 32 host threads (Intel Xeon Gold 6342 at 2.8Ghz). The speedup in MuchiSim’s parallelization is linear until the point where each host thread only processes a couple of tile columns (32-thread 32x32 DUT). fig. 3 demonstrates the efficiency of MuchiSim and its parallelization. On a 32-thread host, simulating the execution of applications on the DUT takes only 43×43$\times$43 × (on geomean, and down to 21×21$\times$21 × for FFT) longer than the expected runtime of every tile of the DUT. This ratio is remarkable considering that MuchiSim performs a functional simulation of data-dependent workloads, where all NoC routers are evaluated every cycle at the flit level.
@@
Figure 4: Simulation time (in host seconds) and throughput in DUT operations and NoC message fits routed per second (y-axis), for scaling DUT sizes from a thousand to a million tiles (x-axis) when processing the RMAT-26 dataset. (This evaluation models 32x32 tiles per chiplet, connected via a 64-bit hierarchical 2D torus.) The $2^{10}$ and $2^{12}$ datapoints are evaluated with 16 and 32 host threads, respectively, of a single-socket Intel Xeon Gold 6342 at 2.8Ghz, the datapoints from $2^{14}$ to $2^{18}$ with 64 and $2^{20}$ with 128 threads, respectively, of a 4-socket Intel Xeon Gold 6230 at 2.1GHz.
@@
fig. 4 shows the simulation time and throughput of MuchiSim (in host seconds) when using increasingly large DUT configurations (from a thousand to a million tiles) to execute the RMAT-26 dataset. The first three datapoints see a decreased or nearly constant simulation time because we are increasing the parallelization of MuchiSim from 16 to 64 host threads. The NoC simulation throughput ranges from a few million message flits routed per second (PageRank) to over 40 million (SSSP). This NoC throughput does not count the routing attempts that fail due to contention (two input ports compete for the same output) or backpressure (the buffer of the output port is full). The throughput of operations—natively executed on the host—reaches up to a few billion Ops/s.
@@
The absolute runtimes in fig. 4 (blue) showcase that even million-tile DUT evaluations on a billion-element dataset—unattainable before MuchiSim—are simulated in under half a day for BFS, SPMV and Histo, and in up to two days for SSSP and PageRank, on a single host server.
FFT is not shown in fig. 4 since for this benchmark we scale the problem size with the system size (i.e., the FFT of a $n^{3}$ tensor is parallelized across $n^{2}$ processors) but we mention some runtimes here. FFT is simulated in under 100 seconds for up to $n=128$ (i.e, $2^{14}$ tiles) and it takes around a day to simulate $n=1024$ (a million tiles on a billion-element tensor), where the all-to-all communication of FFT dominates the runtime.
@@ IV-C Case Study: Memory Integration
Figure 5: Performance, energy efficiency, and performance per dollar improvements of the DUT using different SRAM sizes and number of tiles per HBM channel, over a baseline of 64 KiB SRAM and 128 tiles per HBM channel (Tile/Ch). In this study, a chiplet is always attached to a single 8-channel HBM device, and thus, the number of tiles per chiplet (16x16 or 32x32) determines the ratio of tiles per HBM channel. The RMAT-25 dataset is studied on a DUT with 1024 tiles; the dataset footprint per tile ranges from 4 MiB (Histogram) to 8 MiB per tile (SPMV).
@@
As a case study, we studied the performance, energy efficiency, and performance per dollar improvements of different SRAM sizes and numbers of tiles per chiplet (same number of tiles in total) and show the results in fig. 5. Since all applications but SPMM have a very low arithmetic intensity with respect to the data movement, they require a large amount of memory bandwidth. The top plot of (fig. 5 shows a strong performance increase with SRAM size, 3.5×3.5$\times$3.5 × on geomean when increasing from 64 KiB to 256 KiB, with the same chiplet configuration of 32x32 tiles, and an additional 2×2$\times$2 × increase when using 16x16-tile chiplets (32 Tile/Ch), for the same SRAM size of 256 KiB. The hit-rate of the data cache (not displayed) increases from 83% to 95% on geomean with the SRAM size increase, but since the effective bandwidth of a tile is $SRAM\_bandwidth\times hit\_rate+DRAM\_bandwidth\times(1-hit\_rate)$, the hit-rate has a larger impact on the bandwidth when more tiles share the same DRAM channel (128 Tile/Ch).
@@
The throughput per dollar is lower with 16x16-tile chiplets for all applications but SSSP and SPMV, due to the extra cost of having four times more HBM devices. The throughput per dollar estimates for SPMM—with over an order of magnitude more arithmetic intensity than the rest of the applications for the dataset evaluated (RMAT-25)—showcases that when targeting performance per dollar, the optimal memory integration depends on the application domain.
V Conclusions
@@
MuchiSim is a simulation framework aimed to explore the design space of scale-out architectures for communication-intensive applications like graph analytics and sparse linear algebra. It supports simulating a class of tiled manycore architectures with different hierarchical organizations, network topologies, and memory integrations. It also supports evaluating various parallelization strategies (do-all and task-based) and communication primitives (e.g., message-passing and reduction trees). As demonstrated via a case study, MuchiSim is well suited to explore the optimal balance between resources dedicated to compute, memory and network, for different application domains and target metrics (e.g., performance, power and cost).
The distinguishing factor of MuchiSim is its scalability in the challenging domain of data-dependent and communication-intensive applications, which requires functional simulation (since the execution depends on the data) as well as modeling the communication cycle by cycle. To demonstrate that, we simulated applications with billion-element datasets parallelized with up to a million processing units and presented the simulation times and throughput metrics. We achieve this scalability by optimizing the design and parallelization of the simulator for the target class of tiled manycore architectures.
@@
The MuchiSim framework is open-source and available at our repository[68], including the simulator, applications, datasets, and visualization tools. It also contains the scripts to reproduce the experiments presented in this paper and other case studies, as a tutorial to allow researchers to explore the design space of scale-out manycore architectures. Future work on the simulator itself includes multi-node MPI parallelization and support for low-diameter cluster interconnects, to enable simulating even larger systems.